an object of mass 0.25kg moves at a height,h above the ground with a speed of 4m/s if its mechanical energy at this height 12J determine the value of h
1
Expert's answer
2019-11-08T07:27:35-0500

Mechanical energy = potential energy + kinetic energy = 12 J

Kinetic energy =

"0.5mv^2=0.5\times0.25\times4^2=2\ J"

Then,

Potential energy = "12-2= 10\ J"

So,


"mgh=10\\or\\0.25\times10\times h=10\\or\\h=4\ m"